OpenStudy (anonymous):

(Continuity) What is the limit of abs(x)/x as x approaches 0 from the right?

OpenStudy (valpey):

What is abs(x)/x when x is a very small positive number?

OpenStudy (anonymous):

Since x approaches from the right IxI=x

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Welcome

OpenStudy (valpey):

The obvious next question is going to be: What is the limit of abs(x)/x as x approaches 0 from the left? To this the same thinking goes into: "What is abs(x)/x when x is a very small (in magnitude) negative number?" The general point is that the limit from the left will be different from the limit from the right. Hence the function abs(x)/x is discontinuous at x=0. (Continuity requires those limits to equal each other).

OpenStudy (anonymous):

\[\ |x| = \left\{\begin{array}{rcc} -x & \text{if} & x <0 \\ x& \text{if} &\geq 0 \end{array} \right. \] and so \[\frac{|x|}{x} = \left\{\begin{array}{rcc} -1 & \text{if} & x<0 \\ 1& \text{if} & x >0 \end{array} \right. \]

OpenStudy (anonymous):



OpenStudy (anonymous):

I understand. Thank you